# Zihan Tan    Suyuan Huang    Guancheng Wan    Wenke Huang    He Li    Mang Ye

### v2

@@ Abstract
Federated Graph Learning (FGL) combines the privacy-preserving capabilities of federated learning (FL) with the strong graph modeling capability of Graph Neural Networks (GNNs). Current research addresses subgraph-FL only from the structural perspective, neglecting the propagation of graph signals on spatial and spectral domains of the structure. From a spatial perspective, subgraph-FL introduces edge disconnections between clients, leading to disruptions in label signals and a degradation in the class knowledge of the global GNN. From a spectral perspective, spectral heterogeneity causes inconsistencies in signal frequencies across subgraphs, which makes local GNNs overfit the local signal propagation schemes. As a result, spectral client drifts occur, undermining global generalizability. To tackle the challenges, we propose a global knowledge repository to mitigate label signal disruption and a frequency alignment to address spectral client drifts. The combination of Spatial and Spectral strategies forms our framework $S^{2}$. Extensive experiments on multiple datasets demonstrate the superiority of $S^{2}$. The code is available at https://github.com/Wonder7racer/S2FGL.git
@@ 1 Introduction
Graph Neural Networks (GNNs) have demonstrated remarkable efficacy in modeling graph-structured data (Wan et al., 2025a; Fang et al., 2025), thereby finding applications across various domains, such as social networks (Fan et al., 2020; Zhang et al., 2022b), epidemiology (Liu et al., 2024), and fraud detection (Wang et al., 2019; Tang et al., 2022). However, in real-world scenarios, graph data is often generated at the edge devices rather than in centralized systems (Zhang et al., 2021a). To address this, Federated Graph Learning (FGL) has emerged (Fu et al., 2022; Liu & Yu, 2022; Tan et al., 2025, 2024; Huang et al., 2022; Wan et al., 2024b, 2025b), leveraging the data privacy-preserving capabilities of Federated Learning (FL) (Huang et al., 2024, 2023b, 2023c, 2022) to enable the efficient distributed training of GNNs (Huang et al., 2024). A significant use case of FGL is subgraph-FL, where each participant possesses a subgraph of the same overarching graph data.
Figure 1: (a) According to existing FGL mainstream research, the range of 5 to 20 is the closest to real-world application scenarios. This range is encountering the toughest label signal disruption challenges. (b) The heat map of Kullback-Leibler divergence of eigenvalue distributions across clients. Inconsistency in graph signal frequency of subgraphs leads to spectral client drifts.
@@
Figure 2: Illustration of the problems. (a) From the spatial perspective, nodes in subgraph-FL lose label signals from originally nearby labeled nodes due to edge loss, namely label signal disruption. Correspondingly, GNNs suffer from poor class knowledge, leading to a deteriorated global GNN. (b) From the spectral perspective, spectral heterogeneity induces inconsistencies in signal frequencies across subgraphs, leading to spectral client drifts in the signal propagation schemes of GNNs and degraded global generalizability.
@@
Although numerous FGL methods have attempted to provide solutions based purely on structure, including identifying structurally similar collaborators (Baek et al., 2023; Xie et al., 2021; Li et al., 2024), enhancing structural knowledge exchange (Tan et al., 2023; Huang et al., 2023a; Tan et al., 2025), and retrieving generic information under structural shifts (Wan et al., 2024a; Tan et al., 2024). These approaches overlooked the propagation of graph signals within the structure. Signal propagation can be analyzed from two perspectives: the spatial and the spectral domain. Specifically, the spatial domain governs the explicit transmission of signals among linked nodes, while the spectral domain characterizes signal diffusion across varying frequency spectra.
@@
From the spatial perspective, due to edge loss, we hypothesize that nodes in subgraph-FL lose label signals from originally nearby labeled nodes. This degradation hampers the ability of GNNs to learn comprehensive semantic knowledge, resulting in poor global performance and reduced generalizability. Correspondingly, we define this phenomenon as Label Signal Disruption (LSD), which naturally exists in subgraph-FL. For verifying, inspired by graph active learning research (Han et al., 2023), we investigate how the Structure Inertia Score (SIS) varies in subgraph-FL. Specifically, SIS evaluates the influence and significance of labels on graphs. In Fig. 1(a), we empirically demonstrate that the SIS initially decreases and then increases as the number of partitions increases. Correspondingly, existing methods fail to address LSD and suffer from poor class knowledge. Based on our analysis, we pose the question: I) How can we address label signal disruption in subgraph-FL?
@@
From the spectral perspective, inconsistencies in signal frequencies across subgraphs of clients induce spectral client drifts in the signal transmission schemes of GNNs, thereby undermining the collaboration. To verify this phenomenon, we examine graph spectra across clients and demonstrate the heterogeneity in Fig. 1(b). It reveals inconsistent eigenvalue distribution across clients. As a result, GNNs learn distinct signal propagation schemes of subgraphs and optimize in different spectral directions, leading to spectral client drifts and degraded generalizability. Based on the above analysis, we pose the question: II) How can we alleviate spectral client drifts under spectral heterogeneity?
@@
To address the issue of label signal disruption in Question I), we propose Node Label Information Reinforcement (NLIR). Specifically, we leverage structurally representative nodes to construct a prototype-based global repository of class knowledge. During training, NLIR calculates the similarity distribution between all representative prototypes with node features, which provides multidimensional class localization of nodes. Furthermore, our strategy injects class knowledge into the local GNN, effectively mitigating the LSD issue.
@@
Considering the spectral client drift posed by spectral heterogeneity in II), we propose Frequency-aware Graph Modeling Alignment (FGMA). Our method utilizes the similarity relationship of the nodes to reconstruct a spectrum that incorporates GNNs adjacency awareness. FGMA then projects the high-frequency and low-frequency components of the features onto this spectrum. Subsequently, by aligning the local projections with the global one, we encourage the GNNs to learn a globally generic frequency processing scheme, thereby mitigating spectral client drift.
@@
  • We identify and empirically reveal the issue of label signal disruption in subgraph-FL. In addition, we reveal the spectral heterogeneity in subgraph scenarios.
  • We design our framework $S^{2}$ including strategy Node Label Information Reinforcement and Frequency-aware Graph Modeling Alignment, effectively addressing the label signal disruption and the challenge of spectral client drifts in subgraph-FL.
  • We conduct extensive experiments on various datasets, validating the superiority of our proposed $S^{2}$.
@@ 2 Related Work
Federated Graph Learning. Federated graph learning leverages the powerful graph modeling capabilities of GNNs along with the privacy-preserving attributes of federated learning, thus gaining increasing attention these days (He et al., 2021a; Fu et al., 2022; Liu & Yu, 2022; Wan et al., 2025b). Current FGL research can generally be categorized into two types: intra-graph FGL and inter-graph FGL. Intra-graph FGL research primarily focuses on subgraph-FL scenarios, where each client participates in the collaboration with a part of the whole graph (Zhang et al., 2021b). Correspondingly, the training targets include missing link prediction (Chen et al., 2021; Baek et al., 2023), node classification (Huang et al., 2023a; Li et al., 2024; Wan et al., 2024a; Zhu et al., 2024), and so on. On the other hand, clients in inter-graph FGL own independent local graph data, such as multiple graphs from different domains (Tan et al., 2023; Xie et al., 2021). In this paper, we focus on subgraph-FL scenarios of intra-graph FGL. We are the first to reveal and address the label signal disruption and spectral heterogeneity challenge among subgraphs, while existing methods inevitably fail due to the lack of targeted solutions.
@@
{mdframed}
[backgroundcolor=cadetblue!10, linewidth=0.8pt, linecolor=cadetblue!80, roundcorner=5pt] Graph Signal Propagation: Graph signal propagation describes how node signals diffuse on graph structures. In the spatial domain, propagation occurs through explicit signal passing along edges. In the spectral domain, propagation is characterized by how signals distribute across different frequency components. Label Signal Disruption: As subgraphs experience edge loss, nodes lose critical label signals containing class knowledge from their formerly adjacent labeled neighbors. Consequently, it limits the ability of GNNs to capture class distinctions accurately, leading to a degraded global model. Spectral Client Drifts: Inconsistencies in signal frequencies on graph spectra across subgraphs lead to diverging signal propagation schemes, causing spectral drift and degrading the generalizability of the global model.
@@ 3 Problem Statement
Notation. Let the graph data be represented as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes with $|\mathcal{V}|=N$ vertices, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the set of edges connecting these nodes. The adjacency matrix is represented by $\mathbf{A}\in\mathbb{R}^{N\times N}$, where $\mathbf{A}_{uv}=1$ indicates the presence of an edge $e_{uv}\in\mathcal{E}$, and $\mathbf{A}_{uv}=0$ otherwise. The Laplacian matrix is given by $\mathbf{L}=\mathbf{D}-\mathbf{A}$, where $\mathbf{D}$ is the degree matrix. The unitary matrix $\mathbf{U}$ is composed of the eigenvectors of $\mathbf{L}$. To distinguish between local and global properties, we introduce the following additional notation: the symbol $i$ represents local properties or entities, whereas $g$ denotes global properties or entities. Additional notation includes $\mathbf{h}$, which denotes the feature vector matrix of the graph $\mathcal{G}$, and $\mathbf{h^{\prime}}$, representing the reconstructed feature matrix that emphasizes local structure. The cosine similarity matrix of $\mathbf{h^{\prime}}$ is denoted by $\mathbf{S^{\prime}}$ and the projections of the feature matrix onto low- and high-frequency eigenvectors are represented by $\mathbf{p}_{l}$ and $\mathbf{p}_{h}$, respectively.
@@
Here, $\alpha\in(0,1)$ is the damping factor, typically set to $0.85$, representing the probability of continuing the random walk. $\mathbf{I}$ is the identity matrix of size $N\times N$, $\mathbf{A}$ is the adjacency matrix of the graph, and $\mathbf{D}$ is the degree matrix with $D_{ii}$ denoting the degree of node $i$.
@@
Here, $P$ is the PPR matrix, and $t\in\{0,1\}^{n}$ is a binary vector indicating the training nodes, where $t_{j}=1$ if node $j$ is part of the training set, and $t_{j}=0$ otherwise. The SIS aggregates the maximum personalized PageRank values from each node to any labeled node, effectively measuring the strongest influence each node in the graph receives from the training set. A higher SIS indicates greater structural inertia, suggesting that the labels have a significant influence over the network overall structure.
Figure 3: Framework illustration. (a) Node Label Information Reinforcement (NLIR) leverages a prototype-based structurally representative global repository of class knowledge. It provides multidimensional class localization of nodes through similarity distribution and allows $L_{\textbf{FKD}}$ to inject the class knowledge during training. (b) Frequency-aware Graph Modeling Alignment (FGMA) aligns local high and low spectral adjacency awareness with the global GNN for a generic signal propagation scheme, mitigating spectral client drifts.
@@ 4.1 Motivation
Signal propagation over graph structures fundamentally shapes the signal transmission paradigm in GNNs. Therefore, rather than focusing solely on challenges arising from static graph structures in FGL, as prior works have done, it is crucial to consider the dynamics of signal propagation. Specifically, the spatial domain governs the explicit transmission of signals between connected nodes, while the spectral domain captures signal passing across different frequency components. Accordingly, we empirically validate the presence of two major challenges: label signal disruption and spectral client drifts, from both the spatial and spectral perspectives.
@@
Motivation of NLIR. Graph data is fragmented across multiple clients in subgraph-FL, which inevitably disrupts these label signals, especially as the client scale increases. Label signal disruption severs key pathways for propagating label information, resulting in incomplete and biased local feature representations that ultimately degrade GNNs ability semantically. For validation, we investigate how the Structure Inertia Score (SIS) varies in subgraph-FL and experimentally demonstrate the relationship between the decrease in SIS and the client scale. Specifically, the SIS first decreases and then increases as the number of client partitions increases, highlighting the significant role of LSD in the observed performance decline. We aim to mitigate the impact of label signal disruption in FGL by preserving valuable label knowledge across fragmented subgraphs. Correspondingly, Node Label Information Reinforcement (NLIR) is introduced. By selecting nodes with both structural representativeness and rich label information for a global repository and injecting it during local training, NLIR reduces the information loss inherent in the subgraph scenarios. Specifically, throughout the training process, NLIR assesses the similarity distributions between node features and all representative prototypes for both local and global, thereby enabling multidimensional localization of node classes. By aligning the two similarity distributions, it effectively reconstructs label signal pathways within each client, enhancing local feature modeling by restoring context around labeled nodes.
@@
Motivation of FGMA. We reveal the challenge of spectral client drifts in the subgraph scenario, where GNNs across different clients capture high and low-frequency information inconsistently due to graph structural heterogeneity. This results from overfitting local signal propagation frequencies, which bring biases during collaboration and compromise the generalizability of the global model. To address spectral client drifts, we propose the Frequency-aware Graph Modeling Alignment. Specifically, FGMA reconstructs the local graph spectra by leveraging the GNN adjacency awareness after calculating the node similarity matrix. By projecting features separately onto high- and low-frequency components of the reconstructed spectra and aligh the local and global projections, FGMA promotes the learning of a generalizable spectral signal propagation paradigm, thereby reducing frequency-based discrepancies during collaboration and enhancing the global generalizability.

### v3

@@ Abstract
Federated Graph Learning (FGL) combines the privacy-preserving capabilities of federated learning (FL) with the strong graph modeling capability of Graph Neural Networks (GNNs). Current research addresses subgraph-FL from the structural perspective, neglecting the propagation of graph signals on spatial and spectral domains of the structure. From a spatial perspective, subgraph-FL introduces edge disconnections between clients, leading to disruptions in label signals and a degradation in the semantic knowledge of the global GNN. From a spectral perspective, spectral heterogeneity causes inconsistencies in signal frequencies across subgraphs, which makes local GNNs overfit the local signal propagation schemes. As a result, spectral client drift occurs, undermining global generalizability. To tackle the challenges, we propose a global knowledge repository to mitigate the challenge of poor semantic knowledge caused by label signal disruption. Furthermore, we design a frequency alignment to address spectral client drift. The combination of Spatial and Spectral strategies forms our framework $S^{2}$FGL. Extensive experiments on multiple datasets demonstrate the superiority of $S^{2}$FGL. The code is available at https://github.com/Wonder7racer/S2FGL.git
@@ 1 Introduction
Graph Neural Networks (GNNs) have demonstrated remarkable efficacy in modeling graph-structured data (Wan et al., 2025a; Fang et al., 2025), thereby finding applications across various domains, such as social networks (Fan et al., 2020; Zhang et al., 2022b), epidemiology (Liu et al., 2024), and fraud detection (Wang et al., 2019; Tang et al., 2022). However, in real-world scenarios, graph data is often generated at the edge devices rather than in centralized systems (Zhang et al., 2021a). To address this, Federated Graph Learning (FGL) has emerged (Fu et al., 2022; Liu & Yu, 2022; Tan et al., 2025, 2024; Huang et al., 2022; Wan et al., 2024b, 2025b), leveraging the data privacy-preserving capabilities of Federated Learning (FL) (Huang et al., 2024, 2023b, 2023c, 2022) to enable the efficient distributed training of GNNs (Huang et al., 2024). A prominent application of FGL is subgraph-FL, in which each participant holds a subgraph derived from the same overall graph data.
Figure 1: In the first place, compared with centralized GNN training, subgraph-FL is encountering label signal disruption challenges, leading to decreased structure inertia score and poor semantic knowledge for GNNs. Moreover, we demonstrate the heat map of Kullback-Leibler divergence of eigenvalue distributions across clients. Inconsistency in subgraph signal frequency caused by spectral heterogeneity leads to spectral client drift.
@@
Figure 2: Illustration of the problems. (a) From the spatial perspective, nodes in subgraph-FL lose label signals from originally nearby labeled nodes due to edge loss, namely label signal disruption. Correspondingly, GNNs suffer from poor semantic knowledge, leading to a deteriorated global GNN. (b) From the spectral perspective, spectral heterogeneity induces inconsistencies in signal frequencies across subgraphs, leading to spectral client drift in the signal propagation paradigms of GNNs and degraded global generalizability.
@@
Although numerous FGL methods have attempted to provide solutions based on structure to enhance effectiveness, including identifying structurally similar collaborators (Baek et al., 2023; Xie et al., 2021; Li et al., 2024), enhancing structural knowledge exchange (Tan et al., 2023; Huang et al., 2023a; Tan et al., 2025), and retrieving generic information under structural shifts (Wan et al., 2024a; Tan et al., 2024). Nevertheless, these approaches overlooked the propagation of graph signals within the structure. Specifically, graph signal propagation can be analyzed from two perspectives: the spatial and the spectral domain. Specifically, the spatial domain governs the explicit transmission of signals among linked nodes, while the spectral domain characterizes signal diffusion across varying frequency spectra.
@@
From the spatial perspective, due to edge loss, we hypothesize that nodes in subgraph-FL lose label signals from originally nearby labeled nodes. This degradation hampers the ability of GNNs to learn comprehensive semantic knowledge, resulting in poor global performance and reduced generalizability. Correspondingly, we define this phenomenon as label signal disruption, which naturally exists in subgraph-FL. For verifying, inspired by graph active learning research (Han et al., 2023), we investigate how the Structure Inertia Score (SIS) varies in subgraph-FL. Specifically, SIS evaluates the influence and significance of labels on graphs. In Fig. 1, we empirically demonstrate that the SIS decreases in subgraph-FL compared with centralized training. Correspondingly, existing methods suffer from poor semantic knowledge. Based on our empirical analysis, we pose the question: I) How can we address the challenge of poor semantic knowledge under label signal disruption?
@@
From the spectral perspective, inconsistencies in signal frequencies across clients caused by spectral heterogeneity induce spectral client drift in the signal transmission schemes of GNNs, thereby undermining the collaboration. To verify this phenomenon, we examine graph spectra across clients and demonstrate the heterogeneity in Fig. 1. It reveals inconsistent eigenvalue distribution across clients. As a result, GNNs learn distinct signal propagation schemes of subgraphs and optimize in different spectral directions, leading to spectral client drift and degraded generalizability. Based on our analysis, we pose the question: II) How can we alleviate spectral client drift under spectral heterogeneity?
@@
To address the challenge of poor semantic knowledge under label signal disruption in Question I), we propose Node Label Information Reinforcement (NLIR). Specifically, our strategy leverages structurally and semantically representative nodes to construct a prototype-based global repository of semantic knowledge. During training, NLIR calculates the similarity distribution between all representative prototypes with node features, which provides multidimensional semantic localization of nodes. Consequently, our strategy injects semantic knowledge from the repository into the local GNN during training, effectively mitigating the issue of poor semantic knowledge under label signal disruption.
@@
Considering the spectral client drift posed by spectral heterogeneity in II), we propose Frequency-aware Graph Modeling Alignment (FGMA). Our method utilizes the similarity relationship of the node feature of the frozen global GNN and the local GNN to reconstruct spectra that incorporates GNNs adjacency awareness. FGMA then projects the high-frequency and low-frequency components of the features onto this spectrum. Subsequently, by aligning the local projections with the global one, we encourage the GNNs to learn a globally generic frequency processing scheme, thereby mitigating spectral client drift.
@@
  • First, we identify and empirically reveal the issue of poor semantic knowledge under label signal disruption. In addition, we reveal the spectral client drift under spectral heterogeneity in subgraph-FL.
  • We design our framework $S^{2}$FGL including strategy Node Label Information Reinforcement and Frequency-aware Graph Modeling Alignment, effectively addressing the challenges of poor semantic knowledge and spectral client drift in subgraph-FL.
  • We conduct extensive experiments on various datasets, validating the superiority of our proposed $S^{2}$FGL.
@@ 2 Related Work
Federated Graph Learning. Federated graph learning leverages the powerful graph modeling capabilities of GNNs along with the privacy-preserving attributes of federated learning, thus gaining increasing attention these days (He et al., 2021a; Fu et al., 2022; Liu & Yu, 2022; Wan et al., 2025b). Current FGL research can generally be categorized into two types: intra-graph FGL and inter-graph FGL. Intra-graph FGL research primarily focuses on subgraph-FL scenarios, where each client participates in the collaboration with a part of the whole graph (Zhang et al., 2021b). Correspondingly, the training targets include missing link prediction (Chen et al., 2021; Baek et al., 2023), node classification (Huang et al., 2023a; Li et al., 2024; Wan et al., 2024a; Zhu et al., 2024), and so on. On the other hand, clients in inter-graph FGL own independent local graph data, such as multiple graphs from different domains (Tan et al., 2023; Xie et al., 2021). In this paper, we focus on subgraph-FL scenarios of intra-graph FGL. Specifically, we are the first to empirically reveal and address the challenge of poor semantic knowledge under label signal disruption and client drift under spectral heterogeneity among subgraphs, while existing methods inevitably fail spatially and spectrally due to the lack of targeted solutions.
@@
Graph Signal Propagation: Graph signal propagation describes how node signals diffuse on graph structures. In the spatial domain, propagation occurs through explicit signal passing along edges. In the spectral domain, propagation is characterized by how signals distribute across different frequency components. Label Signal Disruption: As subgraphs experience edge loss, nodes lose critical label signals containing class knowledge from their formerly adjacent labeled neighbors. Consequently, it limits the ability of GNNs to capture class distinctions accurately, leading to poor semantic knowledge under label signal disruption. Spectral Client Drift: Inconsistencies in signal frequencies on graph spectra across subgraphs lead to spectral heterogeneity and diverging signal propagation schemes, causing spectral client drift and degrading the generalizability of the global model.
Figure 3: Framework illustration. (a) Node Label Information Reinforcement (NLIR) leverages a structurally and semantically representative global prototype repository. It provides multidimensional semantic localization of nodes through similarity distribution and allows $L_{\textbf{FKD}}$ to inject the semantic knowledge during training. (b) Frequency-aware Graph Modeling Alignment (FGMA) aligns local high and low spectral adjacency awareness with the global GNN for a generic signal propagation scheme, mitigating spectral drifts.
@@ 3 Problem Statement
Notation. Let the graph data be represented as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is the set of nodes with $|\mathcal{V}|=N$ vertices, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denotes the set of edges connecting these nodes. The adjacency matrix is represented by $\mathbf{A}\in\mathbb{R}^{N\times N}$, where $\mathbf{A}_{uv}=1$ indicates the presence of an edge $e_{uv}\in\mathcal{E}$, and $\mathbf{A}_{uv}=0$ otherwise. Moreover, $\mathbf{X}$ the feature vector matrix of the graph $\mathcal{G}$. The Laplacian matrix is given by $\mathbf{L}=\mathbf{D}-\mathbf{A}$, where $\mathbf{D}$ is the degree matrix. The unitary matrix $\mathbf{U}$ is composed of the eigenvectors of $\mathbf{L}$. To distinguish between local and global properties, we introduce the following notation: the symbol $i$ represents local properties or entities, whereas $g$ denotes global properties or entities.
@@
Here, $\alpha\in(0,1)$ is the teleportation probability, representing the probability of the random walk restarting from the source node. A typical value is $0.15$, which makes the continuation probability $(1-\alpha)=0.85$. $I$ is the identity matrix, $A$ is the adjacency matrix of the graph, and $D$ is the degree matrix with $D_{ii}$ denoting the degree of node $i$.
@@
Here, $P$ is the PPR matrix, and $t\in\{0,1\}^{n}$ is a binary vector indicating the training nodes, where $t_{j}=1$ if node $j$ is part of the training set, and $t_{j}=0$ otherwise. The SIS aggregates the maximum personalized PageRank values from each node to any labeled node, effectively measuring the strongest influence each node in the graph receives from the training set. A higher SIS indicates greater structural inertia, suggesting that the labels have a significant influence over the network overall graph structure.
@@ 4.1 Motivation
Signal propagation over graph structures fundamentally shapes the signal transmission paradigm of GNNs. Therefore, rather than focusing solely on challenges arising from static graph structures in subgraph-FL, it is crucial to consider the dynamics of signal propagation. Correspondingly, we conduct our analysis from both the spatial and spectral domains from the perspective of the graph signal. Specifically, the spatial domain governs the explicit passing of signals between connected nodes, while the spectral domain captures signal diffusion across different frequencies. Accordingly, we empirically validate the presence of two major challenges from the spatial and spectral perspectives: label signal disruption and spectral client drift. These phenomena respectively pose challenges of poor semantic knowledge and spectral client drift, which severely constrain the potential of subgraph-FL collaboration.
@@
Motivation of NLIR. Graph data is fragmented across multiple clients in subgraph-FL, which inevitably disrupts semantic signals from labeled nodes across clients. Label signal disruption leads to key pathways for propagating semantic information, resulting in biased local feature representations that ultimately degrade GNN ability. For validation, we investigate the relationship between the decrease in SIS and the client scale. Specifically, the SIS exhibits a downward trend as the client scale increases, with notably lower scores in the range emphasized by mainstream subgraph-FL studies, thereby highlighting the label signal disruption phenomenon. We aim to mitigate its negative impact by preserving valuable semantic knowledge across fragmented subgraphs. Correspondingly, we propose Node Label Information Reinforcement (NLIR). By selecting nodes with both structural representativeness and rich semantic information for the construction of a global repository and injecting it during local training, NLIR reduces the information loss inherent in the subgraph scenarios. Subsequently, NLIR assesses the similarity distributions between node features and all representative prototypes for both local and global GNNs during local training, thereby enabling multidimensional semantic localization of nodes. By aligning the two similarity distributions, it effectively injects semantic knowledge and enhances feature modeling semantically.
@@
Motivation of FGMA. We reveal the challenge of spectral client drift in subgraph-FL in Fig. 1, where GNNs across different clients capture frequency information inconsistently due to graph spectral heterogeneity. This further leads to overfitting local signal propagation frequencies, which brings spectral conflicts during collaboration and compromises the generalizability of the global GNN. To address spectral client drift, we propose the Frequency-aware Graph Modeling Alignment. Specifically, FGMA reconstructs the local graph spectra with the GNN adjacency awareness by calculating the node similarity matrix. Subsequently, by projecting feature representations separately onto high- and low-frequency components of the reconstructed spectra and aligning the local and global projections, FGMA promotes the learning of a generalizable spectral signal propagation paradigm across clients, thereby reducing frequency-based discrepancies during collaboration and mitigating spectral client drift. Consequently, our strategy effectively enhances the global generalizability spectrally.